A car with wheels of radius 14 in. is driven at 35 mi/h, the wheels turn at an approximate rate of 420 rpm. Show how to obtain this rate of turn.
Answer by Alan3354(69443)

A car with wheels of radius 14 in. is driven at 35 mi/h, the wheels turn at an approximate rate of 420 rpm. Show how to obtain this rate of turn.
-----------
420 rpm = 420*60 = 25200 rev/hour
Circumference = 2*pi*14 inches = 28pi inches
25200*28pi/(12*5280) = mi/hr
= 34.98 mi/hr